# SLO-aware GPU Frequency Scaling for Energy Efficient LLM Inference Serving

### v1

@@ Abstract
As Large Language Models (LLMs) gain traction, their reliance on power-hungry GPUs places ever-increasing energy demands, raising environmental and monetary concerns. Inference dominates LLM workloads, presenting a critical challenge for providers: minimizing energy costs under Service-Level Objectives (SLOs) that ensure optimal user experience. In this paper, we present throttLL’eM, a framework that reduces energy consumption while meeting SLOs through the use of instance and GPU frequency scaling. throttLL’eM features mechanisms that project future KV cache usage and batch size. Leveraging a Machine-Learning (ML) model that receives these projections as inputs, throttLL’eM manages performance at the iteration level to satisfy SLOs with reduced frequencies and instance sizes. We show that the proposed ML model achieves $R^{2}$ scores greater than 0.97 and miss-predicts performance by less than 1 iteration per second on average. Experimental results on LLM inference traces show that throttLL’eM achieves up to 43.8% lower energy consumption and an energy efficiency improvement of at least 1.71×1.71$\times$1.71 × under SLOs, when compared to NVIDIA’s Triton server.
I Introduction
@@
Large Language Models (LLMs) have surged in popularity over the last couple of years, driven by advancements in deep learning and highlighted by the remarkable success of ChatGPT from OpenAI [3]. This rapid adoption, combined with the reliance of LLMs on high-end GPUs with significant power and energy consumption has raised concerns about their environmental impact. [48, 41]. Recent studies indicate that inference dominates the computational workload, accounting for over 90% of the overall LLM compute cycles within data centers[44, 42]. Although individual inference tasks are short-lived, the energy required for a single response from an LLM like GPT-4 can be significant, reaching $\approx 3$Wh [7]. As userbases for LLM services grow into the millions [1], their energy footprint is projected to become substantial, potentially reaching 1050 TWh by 2026 [22], rivaling entire countries.
@@
Minimizing the energy footprint of LLM inference serving systems is crucial not only for environmental reasons but also from the providers’ perspective. Efficient utilization of resources allows data center providers to reduce their operating costs, while also facilitating compliance to power budget constraints, enforced by contracts with utility companies [67]. However, reducing energy consumption often conflicts with the need for high performance. LLM-based services (e.g., ChatGPT) are highly interactive, as users submit their queries (e.g., questions) and await for “on-the-fly” responses. This user expectation of low latency is often quantified through Service-Level-Objectives (SLO), e.g., maintaining the $99^{th}$ percentile of response latency under a specific threshold [43].
@@
    Autoregressive nature of LLMs: Unlike traditional feedforward neural networks (e.g., CNNs) with fixed execution steps, LLMs exhibit inherent stochasticity in their inference process. Response generation proceeds iteratively, processing one token (word or sub-word unit) at a time until an End-of-Sequence (EOS) token is generated or a predefined maximum token limit is reached. This dynamic token generation length makes it difficult to precisely pre-allocate computational resources for each inference request.
@@
    Dynamic Batch Composition: Techniques like inflight batching [65] leverage the autoregressive nature of LLMs to improve throughput by allowing requests to enter and exit a batch dynamically. However, as shown in Sec. III-A1, different batch sizes introduce significant performance variability, possibly increasing time-between-tokens and end-to-end latency by up to 45%, thus complicating energy optimization under SLOs.
@@
    Variable Memory Footprint: LLMs exhibit variable memory footprint due to their autoregressive nature. Within each iteration, memory is dynamically reserved for the growing sequence length and only deallocated once a request is completed. This takes the form of a variably sized Key-Value (KV) cache, which stores the information associated with each token. While techniques like paged attention [28] have mitigated memory fragmentation issues, managing the KV cache remains challenging. Notably, increased KV cache usage can lead to an up to 18.2% performance degradation in the inference process (cf. Sec. III-B).
Managing the energy consumption of inference serving systems through traditional power-capping [31] and power-oversubscription [41, 27] techniques falls short in the case of LLMs. These solutions often operate in a static manner, leading to high tail latencies under load [68], potentially conflicting with the specified SLOs set by providers [48]. Moreover, techniques designed for energy optimization of traditional neural network inference serving systems [64], such as delayed [14] or coordinated[37] batching, cannot be directly applied to LLMs due to their unpredictability. Finally, race-to-idle strategies [26, 29] are also rendered inefficient since LLM inference serving systems typically experience a continuous stream of requests [43, 59], making it challenging to achieve true idle states for extended durations.
@@
However, effectively applying frequency scaling to LLM inference remains challenging due to the inherent unpredictability in the text generation process exhibited by these models. While recent works address generation length prediction [24, 46, 21, 71], there is a notable gap on modeling the impact of LLM-intrinsic characteristics and leveraging these insights for energy-efficient LLM inference serving.
@@ II Background on LLM Inference
Figure 1: LLM’s a) serving stack b) text generation process.
LLMs are neural networks with billions of parameters that heavily rely on the Transformer architecture [56]. Figure 1a shows an overview of a typical LLM inference serving stack. Inference servers (e.g., Triton [40], LMDeploy [13]) bridge the gap between users and pre-trained LLMs. They act as a frontend, receiving user queries and interacting with inference serving backends (e.g., TensorRT-LLM backend [39], vLLM [28]), which handle the deployment of LLM instances on the underlying infrastructure.
@@
LLM text generation process: LLMs generate text in a step-by-step manner, one token (word or subword) at a time, as shown in Fig 1b. The process begins with a prompt from the user, which is tokenized before being fed into the model. The tokenization vector is passed to the LLM, generating the first token. This is known as the prefill or prompt phase and is compute-bound [43]. The LLM then enters an iterative loop until it generates an End-of-Sequence (EOS) token or reaches a maximum length limit (max_tokens). During each iteration, the LLM considers the prompt and previously generated tokens to calculate token occurrence probabilities and select a new one. The generated token is appended to the sequence, and the process repeats. This phase is referred to as decoding or generation and is known to be memory bound [4].
@@
Inflight Batching: Inflight batching [65] is an optimization technique which capitalizes on the autoregressive nature of LLMs, by allowing new requests to be added/removed to/from an existing batch while it is being processed.
Instance Parallelism: Modern LLM serving frameworks [39, 28] support partitioning (sharding), enabling model distribution among many GPUs. This allows for executing large models on GPUs that would otherwise not be able to store LLM weights individually. Three key approaches exist to distribute the workload across multiple GPUs [38]. Distributed Data parallelism (DDP) creates identical copies of the model across multiple GPUs. Tensor parallelism (TP) splits the model’s weight tensors across multiple GPUs, enabling parallel computation for specific operations. Finally, Pipeline parallelism (PP) assigns consecutive LLM layers or segments to different GPUs, similar to instruction pipelining in CPUs.
LLM inference performance metrics: Several metrics have been proposed to evaluate LLM performance and serve as potential SLOs[43, 54]. Time To First Token (TTFT) measures the responsiveness of the inference engine, i.e., the time to generate the first token after query submission. Throughput metrics include Time Between Tokens (TBT), LLM iterations per second (IPS), and tokens per second (TPS). End To End latency (E2E) represents the total time from query submission to completion. To quantify energy efficiency, we consider tokens per Joule (TPJ), measuring the number of tokens generated per unit of energy expended.
III Energy Efficient LLM Inference Serving: Challenges & Opportunities
In this section we examine the impact of various factors that affect LLM performance and energy efficiency. We perform this analysis on a NVIDIA DGX A100 system, utilizing the Llama2-13B model as a representative example of LLMs  [^1] Since LLMs typically consist of continuous Attention blocks, similar conclusions can be drawn for other models., with a tensor parallelism of two, unless otherwise stated.
III-A Performance-Energy Trade-offs in LLM Inference
@@
III-A1 Impact of Batch Size and GPU Frequency
We first examine the impact of batch size and GPU frequency on the performance and power consumption of LLM inference. To this end, we spawn batches of different sizes, ranging from 1 to 32 requests wide. To avoid any variation or bias due to the context of the queries, all queries are identical with their prompt and generation lengths fixed to a specific value of 1 input token and 1024 generation tokens. Figures 2a-2e show the average throughput, E2E latency, TBT, power consumption and energy efficiency for the lifetime of each batch.
@@ III-A1 Impact of Batch Size and GPU Frequency
Energy Efficiency Analysis: To further support the above statement, Fig. 2e also presents the energy efficiency of the system, i.e., the number of tokens generated per Joule of energy consumed. We see that processing larger batches improves energy efficiency for all frequencies. Counterintuitively to Fig. 2d, reductions in frequency do not always correspond to increased efficiency. We observe a clear sweet spot: efficiency increases with lower GPU frequencies up to a point (1050 MHz) before dropping again. Notably, using the maximum batch size with a lower, more efficient 1050 MHz frequency can yield a significant (37.4%) boost in energy efficiency with minimal impact on performance (-6.25%, +8.26%, +5.41% throughput, E2E latency, and TBT respectively). For low frequencies ($<$840 MHz) the reduction in power is offset by lower throughput, diminishing overall efficiency while simultaneously worsening performance. The above observations suggest that frequency scaling offers a valuable control knob for energy-efficient LLM serving. By dynamically adjusting the GPU frequency based on workload size (batch size) and aiming for the sweet spot where efficiency is maximized, we can achieve significant energy savings without sacrificing substantial performance. We note that optimal GPU frequency is architecture- and LLM-dependent and may vary with different generation lengths.
@@
III-B Implications of KV cache usage on LLM inference
@@
III-C LLM partitioning implications on performance and energy
@@
Figure 4: Impact of different LLM partitioning approaches.
As described in Sec. II, LLMs support various techniques for partitioning the model across multiple GPUs, each impacting system efficiency differently. Figures 4a and 4b show the throughput (TPS) and energy efficiency (TPJ) respectively, for different partitioning techniques (DDP, PP, TP) and parallelism levels across various batch sizes. Regarding throughput, TP consistently outperforms DDP/PP, exhibiting a throughput improvement of 1.54/2.74×1.54/2.74$\times$1.54 / 2.74 × and 1.79/6.26×1.79/6.26$\times$1.79 / 6.26 × for parallelism levels 2 and 4 respectively, at the maximum batch size supported by all configurations. Notably, TP also supports larger attainable batch sizes, demonstrating its ability to handle more concurrent requests and provide increased serving capacity for an equal number of GPUs. As expected, increasing the level of parallelism enhances performance due to the increased computational resources and memory capacity provided by additional GPUs, which results in more available KV cache blocks, allowing the system to process a greater number of requests simultaneously. Efficiency-wise, TP again outperforms DDP and PP. However, smaller engine sizes provide better energy efficiency up to the maximum batch size they can support. Specifically, we see that TP2 achieves up to 9.66% higher TPJ compared to TP4, when running close to its maximum batch size, indicating that smaller engines should be prioritized whenever possible to optimize energy efficiency.
$\star$ Takeaways: TP configuration outperforms DDP and PP in both performance and energy efficiency. The trade-off between performance and energy efficiency at different parallelism levels underscores the need for dynamic engine sizing based on real-time workload characteristics.
III-D Unpredictability in LLM inference serving systems
Inference serving systems exhibit significant unpredictability due to the diversity of input prompts and workload traffic variations [43, 59]. To illustrate this, we analyze a 60-minute production trace from Azure LLM inference services [43].
@@ III-D Unpredictability in LLM inference serving systems
Figure 5: Analysis of Azure’s LLM inference trace [43].
Length Distribution: Figure 5a shows the distribution of prompt tokens (top) and generated tokens (bottom). Query prompts can reach up to 4000 tokens, with most queries falling between 0-1500 tokens. Similarly, generated tokens range from 10 to 700, with the majority between 100 and 400 tokens. The long-tail distributions observed for both prompt and generated tokens underscore the necessity for dynamic resource allocation strategies to handle the varying computational demands of different queries. Moreover, the unpredictable nature of token generation lengths further complicates resource management due to its direct impact on KV cache utilization and overall system performance (as shown in Sec. III-B).
@@
IV throttLL’eM: SLO-aware & Energy-Efficient LLM Inference Serving
Building on insights from Sec. III, we introduce throttLL’eM, an energy-efficient, SLO-aware LLM inference serving system. throttLL’eM prioritizes critical SLOs, namely E2E and TBT, delivering energy-efficient inference by dynamically adjusting GPU frequency to match workload demands. Figure 6 shows an overview of throttLL’eM. New queries are processed by the generation length prediction component, which estimates the number of generated tokens. This prediction enables accurate KV cache usage and batch size projections, allowing the system to evaluate potential SLO violations. If no violations are anticipated, the query is scheduled; otherwise, it is queued. Finally, throttLL’eM also includes an autoscaling mechanism to dynamically adjust LLM instance parallelism based on workload, and a throttling controller to optimize GPU frequency and energy consumption. Table I summarizes the terminology used in this section.
TABLE I: Terminology used in Section IV
@@
$KV$ Vector of predicted KV cache usage for the LLM engine
@@
$N$ Number of tokens per KV cache block
@@
IV-A Generation Length Prediction
Accurate generation length prediction is essential for optimizing LLM inference, as it can mitigate KV cache overheads (cf. Sec. III-B) and prevent out-of-memory errors [24] due to dynamic memory allocations [28]. Given the direct correlation between the number of generated tokens and allocated KV blocks, length prediction is crucial for preventing such issues. However, accurately predicting generation length is challenging due to the stochasticity of LLMs, the diversity of input queries, and other factors influencing output length.
Recent scientific works have proposed various approaches to tackle the aforementioned problem [24, 46, 21, 71]. In [24, 21], the authors model length prediction as a classification task, fine-tuning DistilBERT [49] and OPT-125M [69] respectively, to classify output length into “buckets”. In contrast, the authors of [46] use a regression model, appending a two-layer fully connected network to a BERT base model[17]. In [71], the authors propose the ”Perception in Advance” method, where the model estimates the length of its response before generating the output. All approaches report errors up to $\approx 30$%. In the context of throttLL’eM, we consider a reliable length predictor that can accurately estimate the expected generation length ($\widehat{|r_{a}|}$) of each input query. This predicted token generation length ($\widehat{|r_{a}|}$) is propagated to the KV usage & Batch Size projection component.

### v2

# throttLL’eM: Predictive GPU Throttling for Energy Efficient LLM Inference Serving

@@ Abstract
As Large Language Models (LLMs) gain traction, their reliance on power-hungry GPUs places ever-increasing energy demands, raising environmental and monetary concerns. Inference dominates LLM workloads, presenting a critical challenge for providers: minimizing energy costs under Service-Level Objectives (SLOs) that ensure optimal user experience. In this paper, we present throttLL’eM, a framework that reduces energy consumption while meeting SLOs through the use of instance and GPU frequency scaling. throttLL’eM features mechanisms that project future Key-Value (KV) cache usage and batch size. Leveraging a Machine-Learning (ML) model that receives these projections as inputs, throttLL’eM manages performance at the iteration level to satisfy SLOs with reduced frequencies and instance sizes. We show that the proposed ML model achieves $R^{2}$ scores greater than 0.97 and miss-predicts performance by less than 1 iteration per second on average. Experimental results on LLM inference traces show that throttLL’eM achieves up to 43.8% lower energy consumption and an energy efficiency improvement of at least $1.71\times$ under SLOs, when compared to NVIDIA’s Triton server. throttLL’eM is publicly available at https://github.com/WilliamBlaskowicz/throttLL-eM.
1 Introduction
@@
Large Language Models (LLMs) have surged in popularity over the last couple of years, driven by advancements in deep learning and highlighted by the remarkable success of ChatGPT from OpenAI [achiam2023gpt]. This rapid adoption, combined with the reliance of LLMs on high-end GPUs with significant power and energy consumption has raised concerns about their environmental impact. [samsi2023words, patel2023polca]. Recent studies indicate that inference dominates the computational workload, accounting for over 90% of the overall LLM compute cycles within data centers[patterson2022carbon, patel2024characterizing]. While the environmental impact of data centers is primarily attributed to hardware manufacturing [chasingCarbon], their operational footprint is also significant; the energy required for a single GPT-4 response can reach $\approx 3$Wh [energ-per-query]. As userbases for LLM services grow into the millions [chatgpt_stats], their energy footprint is projected to become substantial, potentially reaching 1050 TWh by 2026 [iea-electricity], rivaling entire countries.
@@
Minimizing the energy footprint of LLM inference serving systems is crucial not only for environmental reasons but also from the providers’ perspective. Efficient utilization of resources allows data center providers to reduce their operating costs, while also facilitating compliance to power budget constraints, enforced by contracts with utility companies [zhang2021flex]. However, reducing energy consumption often conflicts with the need for high performance. LLM-based services (e.g., ChatGPT) are highly interactive, as users submit their queries (e.g., questions) and await for “on-the-fly” responses. This user expectation of low latency is often quantified through Service-Level-Objectives (SLO), e.g., maintaining the $99^{th}$ percentile of response latency under a specific threshold [patel2024splitwise].
@@
    Autoregressive nature of LLMs: Unlike traditional feed-forward neural networks (e.g., CNNs) with fixed execution steps, LLMs exhibit inherent stochasticity in their inference process. Response generation proceeds iteratively, processing one token (word or sub-word unit) at a time until an End-of-Sequence (EOS) token is generated or a predefined maximum token limit is reached. This dynamic token generation length makes it difficult to precisely pre-allocate computational resources for each inference request.
@@
    Dynamic Batch Composition: Techniques like inflight batching [gyeong2022orca] leverage the autoregressive nature of LLMs to improve throughput by allowing requests to enter and exit a batch dynamically. However, as shown in Sec. 3.1.1, different batch sizes introduce significant performance variability, possibly increasing time-between-tokens and end-to-end latency by up to 45%, thus complicating energy optimization under SLOs.
@@
    Variable Memory Footprint: LLMs exhibit variable memory footprint due to their autoregressive nature. Within each iteration, memory is dynamically reserved for the growing sequence length and only deallocated once a request is completed. This takes the form of a variably sized Key-Value (KV) cache, which stores the information associated with each token. While techniques like paged attention [kwon2023pagedattention] have mitigated memory fragmentation issues, managing the KV cache remains challenging. Notably, increased KV cache usage can lead to an up to 18.2% performance degradation in the inference process (cf. Sec. 3.2).
Managing the energy consumption of inference serving systems through traditional power-capping [li2020thunderbolt] and power-oversubscription [patel2023polca, kumbhare2021prediction] techniques falls short in the case of LLMs. These solutions often operate in a static manner, leading to high tail latencies under load [MarkInferenceServing], potentially conflicting with the specified SLOs set by providers [samsi2023words]. Moreover, techniques designed for energy optimization of traditional neural network inference serving systems [ye2023deep], such as delayed [crankshaw2017clipper] or coordinated[nabavinejad2022coordinated] batching, cannot be directly applied to LLMs due to their unpredictability. Finally, race-to-idle strategies [kim2015racing, le2011slow] are also rendered inefficient since LLM inference serving systems typically experience a continuous stream of requests [patel2024splitwise, wang2024burstgptrealworldworkloaddataset], making it challenging to achieve true idle states for extended durations.
@@
However, effectively applying frequency scaling to LLM inference remains challenging due to the inherent unpredictability in the text generation process exhibited by these models. While recent works address generation length prediction [jin2023s3, qiu2024efficient, hu2024inference, NEURIPS2023_ce7ff340], there is a notable gap in modeling the impact of LLM-intrinsic characteristics and leveraging these insights for energy-efficient LLM inference serving.
@@ 2 Background on LLM Inference
Figure 1: LLM a) serving stack and b) text generation process.
LLMs are neural networks with billions of parameters that heavily rely on the Transformer architecture [vaswani2017attention]. Figure 1a shows an overview of a typical LLM inference serving stack. Inference servers (e.g., Triton [triton-inference-server], LMDeploy [2023lmdeploy]) bridge the gap between users and pre-trained LLMs. They act as a frontend, receiving user queries and interacting with inference serving backends (e.g., TensorRT-LLM backend [tensorRT-LLM-backend], vLLM [kwon2023pagedattention]), which handle the deployment of LLM instances on the underlying infrastructure.
@@
LLM text generation process: LLMs generate text in a step-by-step manner, one token (word or sub-word unit) at a time, as shown in Fig 1b. The process begins with a prompt from the user, which is tokenized before being fed into the model. The tokenization vector is passed to the LLM, generating the first token. This is known as the prefill or prompt phase and is compute-bound [patel2024splitwise]. The LLM then enters an iterative loop until it generates an End-of-Sequence (EOS) token or reaches a maximum length limit (max_tokens). During each iteration, the LLM considers the prompt and previously generated tokens to calculate token occurrence probabilities and select a new one. The generated token is appended to the sequence, and the process repeats. This phase is referred to as decoding or generation and is known to be memory bound [agrawal2023sarathi].
@@
Inflight Batching: Inflight batching [gyeong2022orca] is an optimization technique which capitalizes on the autoregressive nature of LLMs, by allowing new requests to be added to or removed from an existing batch while it is being processed.
Instance Parallelism: Modern LLM serving frameworks [tensorRT-LLM-backend, kwon2023pagedattention] support partitioning (sharding), enabling model distribution among many GPUs. This allows for executing large models on GPUs that would otherwise not be able to store LLM weights individually. Three key approaches exist to distribute the workload across multiple GPUs [nvidia-parallelisms]. Distributed Data parallelism (DDP) creates identical copies of the model across multiple GPUs. Tensor parallelism (TP) splits the model’s weight tensors across multiple GPUs, enabling parallel computation for specific operations. Finally, Pipeline parallelism (PP) assigns consecutive LLM layers or segments to different GPUs, similar to instruction pipelining in CPUs.
LLM inference performance metrics: Several metrics have been proposed to evaluate LLM performance and serve as potential SLOs[patel2024splitwise, stojkovic2024towards]. Time To First Token (TTFT) measures the responsiveness of the inference engine, i.e., the time to generate the first token after query submission. Throughput metrics include Time Between Tokens (TBT), LLM iterations per second (IPS), and tokens per second (TPS). End To End latency (E2E) represents the total time from query submission to completion. To quantify energy efficiency, we consider tokens per Joule (TPJ), measuring the number of tokens generated per unit of energy expended.
3 Energy Efficient LLM Inference Serving: Challenges & Opportunities
In this section we examine the impact of various factors that affect LLM performance and energy efficiency. We perform this analysis on a NVIDIA DGX A100 system, utilizing the Llama2-13B model as a representative example of LLMs  [^1] Since most LLMs are built upon a backbone of stacked Transformer blocks, similar conclusions can be drawn for other models., with a tensor parallelism of two, unless otherwise stated.
3.1 Performance-Energy Trade-offs in LLM Inference
@@
3.1.1 Impact of Batch Size and GPU Frequency
We first examine the impact of batch size and GPU frequency on the performance and power consumption of LLM inference. To this end, we spawn batches of different sizes, ranging from 1 to 32 requests wide. To avoid any variation or bias due to the context of the queries, all queries are identical with their prompt and generation lengths fixed to a specific value of 1 input token and 1024 generation tokens. Figures 2a-2e show the average throughput, E2E latency, TBT, power consumption, and energy efficiency during the lifetime of each batch.
@@ 3.1.1 Impact of Batch Size and GPU Frequency
Energy Efficiency Analysis: To further support the above statement, Fig. 2e presents the energy efficiency of the system, i.e., the number of tokens generated per Joule. We see that processing larger batches improves energy efficiency for all frequencies. Counterintuitively to Fig. 2d, reductions in frequency do not always correspond to increased efficiency. We observe a clear sweet spot: efficiency increases with lower GPU frequencies up to a point (1050 MHz) before dropping again. Notably, using the maximum batch size with a lower, more efficient 1050 MHz frequency can yield a significant 37.4% boost in energy efficiency with minimal impact on performance (-6.25%, +8.26%, +5.41% throughput, E2E latency, and TBT respectively). For low frequencies (<< 840 MHz) the reduction in power is offset by lower throughput, diminishing overall efficiency while simultaneously worsening performance. The above observations suggest that frequency scaling offers a valuable control knob for energy-efficient LLM serving. By dynamically adjusting the GPU frequency based on workload size (batch size) and aiming for the sweet spot where efficiency is maximized, we can achieve significant energy savings without sacrificing substantial performance. We note that optimal GPU frequency is architecture- and LLM-dependent and may vary with different generation lengths.
@@
3.2 Implications of KV cache usage on LLM inference
@@
3.3 LLM partitioning implications on performance and energy
@@
Figure 4: Impact of different LLM partitioning approaches on a) throughput and b) energy efficiency.
As described in Sec. 2, LLMs support various techniques for partitioning the model across multiple GPUs, each impacting system efficiency differently. Figures 4a and 4b show the throughput (TPS) and energy efficiency (TPJ) respectively, for different partitioning techniques (DDP, PP, TP) and parallelism levels across various batch sizes on a single node. Regarding throughput, TP consistently outperforms DDP/PP, exhibiting a throughput improvement of $1.54/2.74\times$ and $1.79/6.26\times$ for parallelism levels 2 and 4 respectively, at the maximum batch size supported by all configurations. Notably, TP also supports larger attainable batch sizes, demonstrating its ability to handle more concurrent requests and provide increased serving capacity for an equal number of GPUs. As expected, increasing the level of parallelism enhances performance due to the increased computational resources and memory capacity provided by additional GPUs, which results in more available KV cache blocks, allowing the system to process a greater number of requests simultaneously. Efficiency-wise, TP again outperforms DDP and PP. However, smaller engine sizes provide better energy efficiency up to the maximum batch size they can support. Specifically, we see that TP2 achieves up to 9.66% higher TPJ compared to TP4, when running close to its maximum batch size, indicating that smaller engines should be prioritized whenever possible to optimize energy efficiency.
$\star$ Takeaways: For intra-node scaling, TP configuration outperforms DDP and PP in both performance and energy efficiency. The trade-off between performance and energy efficiency at different levels of parallelism underscores the need for dynamic engine sizing based on real-time workload characteristics.
3.4 Unpredictability in LLM inference serving systems
Inference serving systems exhibit significant unpredictability due to the diversity of input prompts and workload traffic variations [patel2024splitwise, wang2024burstgptrealworldworkloaddataset]. To illustrate this, we analyze a 60-minute production trace from Azure LLM inference services [patel2024splitwise].
@@ 3.4 Unpredictability in LLM inference serving systems
Figure 5: Analysis of Azure’s LLM inference trace [patel2024splitwise].
Length Distribution: Figure 5a shows the distribution of prompt tokens (top) and generated tokens (bottom). Query prompts can reach up to 4000 tokens, with most queries falling between 0-1500 tokens. Similarly, generated tokens range from 10 to 700, with the majority between 100 and 400 tokens. The long-tail distributions observed for both prompt and generated tokens underscore the necessity for dynamic resource allocation strategies to handle the varying computational demands of different queries. Moreover, the unpredictable nature of token generation lengths further complicates resource management due to its direct impact on KV cache utilization and overall system performance (as shown in Sec. 3.2).
@@
4 throttLL’eM: SLO-aware & Energy-Efficient LLM Inference Serving
Building on insights from Sec. 3, we introduce throttLL’eM, an energy-efficient, SLO-aware LLM inference serving system. throttLL’eM prioritizes critical SLOs, namely E2E and TBT, delivering energy-efficient inference by dynamically adjusting GPU frequency to match workload demands. Figure 6 shows an overview of throttLL’eM. New queries are processed by the generation length prediction component, which estimates the number of generated tokens. This prediction enables accurate KV cache usage and batch size projections, allowing the system to evaluate potential SLO violations. If no violations are anticipated, the query is scheduled; otherwise, it is queued. Finally, throttLL’eM also includes an autoscaling mechanism to dynamically adjust LLM instance parallelism based on workload, and a throttling controller to optimize GPU frequency and energy consumption. Table 1 summarizes the terminology used in this section.
Table 1: Terminology used in Section 4
@@
KVKV Vector of predicted KV cache usage for the LLM engine
@@
NN Number of tokens per KV cache block
@@
$\max{F}$ The maximum frequency supported by the GPU
$E$ The size of the engine (number of GPUs)
4.1 Generation Length Prediction
Accurate generation length prediction is essential for optimizing LLM inference, as it can mitigate KV cache overheads (cf. Sec. 3.2) and prevent out-of-memory errors [jin2023s3] due to dynamic memory allocations [kwon2023pagedattention]. Given the direct correlation between the number of generated tokens and allocated KV blocks, length prediction is crucial for preventing such issues. However, accurately predicting generation length is challenging due to the stochasticity of LLMs, the diversity of input queries, and other factors influencing output length.
Recent works have proposed various approaches to tackle the aforementioned problem [jin2023s3, qiu2024efficient, hu2024inference, NEURIPS2023_ce7ff340]. In [jin2023s3, hu2024inference], the authors model length prediction as a classification task, fine-tuning DistilBERT [sanh2019distilbert] and OPT-125M [zhang2022opt] respectively, to classify output length into “buckets”. In contrast, the authors of [qiu2024efficient] use a regression model, appending a two-layer fully connected network to a BERT base model[devlin2018bert]. In [NEURIPS2023_ce7ff340], the authors propose the "Perception in Advance" method, where the model estimates the length of its response before generating the output. All approaches report errors up to $\approx 30$%. In this paper, we consider a prediction model that estimates the expected generation length ($\widehat{|r_{a}|}$) of each input query. This predicted token generation length ($\widehat{|r_{a}|}$) is propagated to the KV cache usage & Batch Size projection component.